OpenStudy (anonymous):

A stone is swinging in a horizontal circle of diameter 0.8m at 30 Revolution/min. A distant light causes a shadow of the stone to fall on a nearby wall. The amplitude and time period of the SHM of the shadow are?????

OpenStudy (binary3i):

 amplitude is .4m 30 revolution in one minute for the stone. find the time taken by the stone in one revolution which will also be equal to the time taken by the shadow to complete its one cycle of harmonic motion, its time period. 30/min=30/(60sec)=1/(2sec). this means in one round of stone it takes two seconds. so this is the time period of stone and you also have the time period of shadow. amplitude- .4m Time period- 2sec